OpenStudy (anonymous):

How to solve: 4x - 7y=15 -8 + 14y =-30 Using the elimination method?

OpenStudy (anonymous):

elimination would probably be easiest because if you solve by substitution, you will end up with fractions

OpenStudy (anonymous):

oh...wait....divide everything in the second equation by 2

OpenStudy (anonymous):

but how do I solve it? I don't know how to do elimination.

OpenStudy (anonymous):

-8x + 14y = -30 -- divide by 2 -4x + 7y = -15 now we have : 4x - 7y = 15 -4x + 7y = -15 ---------------add 0 + 0 = 0 There is INFINITE SOLUTIONS to this equation